PV of annuity due

5) find the present value of annuity due of 7,000 payable at the beginning of each quarter for six years if the interest rate is 14% compounded quarterly.

6) what is the cash prize of a freezer than can be bought for 7500 a quarter for 2&1/2 if the interest payment is made now and the interest rate is 7%compounded quarterly?


1
Expert's answer
2021-05-23T19:30:01-0400

5.)

Annuity Due refers to series of equalized payments that are made at START of definite interval for finite time period.

Compute Present Value(PV) as below:




Resultant table:





Hence, PV is $116,342.87


6.)

GIVEN,

A= 7500

N = 2.5

R=7%

M = 4 ( QUARTERLY PAYMENT)


"PV=A\times(\frac{1-(1+\frac{R}{M})^{-M\times N}}{\frac{R}{M}})"


"=7500\times(\frac{1-(1+\frac{0.07}{4})^{-4\times 2.5}}{\frac{0.07}{4}})"


"=7500\times9.10\\=68,259.17"



 ANSWER: 68,259.17